# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v2

@@ Abstract.
Research-paper blog posts help scientists to disseminate their work to a larger audience, but translating scientific long documents into long-form summaries like blog posts raises unique challenges: 1) planning what paper content to include in the blog post, 2) drafting the selected content in sections amenable to a paper blog post, and 3) revising the blog post to be scientifically accurate but also concise, easy to understand, and engaging. Can we harness the power of large language models (LLMs) to assist researchers with these challenges? To investigate this question, we developed Papers-to-Posts, an LLM-powered tool that implements a new Plan-Draft-Revise workflow for mixed-initiative long-form paper summarization. An LLM-generated paper outline with pre-selected yet adjustable bullet points helps users to plan what information to include. Meanwhile, customizable LLM instructions support drafting the text with a suitable structure and revising the text to have an appropriate tone. Through two studies, we compared Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form LLM prompting, and we found that Papers-to-Posts improved researchers’ editing power. In a within-subjects lab study (N=20 participants), Papers-to-Posts led participants to make significantly more change to initial LLM drafts within a fixed amount of time and to be significantly more satisfied with their final blog post, without increasing cognitive load. Furthermore, in a between-subjects deployment study (N=37 blog posts, 26 participants), Papers-to-Posts led participants to make more change to initial LLM drafts within a given amount of time as well as writing actions, without decreasing satisfaction with the final blog posts or increasing cognitive load.
human-AI interaction, generative AI, human-AI summarization, science communication, writing support tools
$c$s: Human-centered computing Interactive systems and toolsccs: Human-centered computing User studiesccs: Human-centered computing Natural language interfaces
Figure 1. The Papers-to-Posts system. a) To start, the LLM generates an outline of the input paper. It then produces a blog post draft in which each section is based on section-specific instructions, section-specific bullet points selected by the LLM, and the sections written before it. Given the draft, the user may follow the Plan-Draft-Revise workflow for each blog post section. b) During the planning stage, the user selects a section on which to work and adjusts the tool’s selection of bullet points in the generated paper outline. c) During the drafting stage, the user customizes the various section-specific LLM instructions to fit their needs. d) During the revising stage, the user utilizes macros for modifying text to have an appropriate style for their blog post. (Example output is in Appendix F.)
@@ 1. Introduction
Sharing scientific ideas outside of research papers is important for increasing awareness and impact (Jucan and Jucan, 2014; Burns et al., 2003). Scientists often seek different ways to communicate their work to broader audiences, such as through tweets or blog posts (Gero et al., 2021; Long et al., 2023; Williams et al., 2022; Holmberg and Thelwall, 2014); indeed, tweeting about research papers has been shown to raise their citation count (Luc et al., 2021). While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create longer-form derivatives like blog posts, which require more extensive planning, drafting, and revising writing processes (Flower and Hayes, 1981). Integrating large language models (LLMs) in writing tools offers a promising avenue for improving editing power (i.e., change in blog post per minute or writing action). However, existing approaches are not designed to help create long-form derivatives like blog posts from scientific long documents like academic papers.
Prior work in human-LLM text summarization (Zhang et al., 2023b; Lai et al., 2022) and human-LLM scientific writing (Long et al., 2023; Kim et al., 2023) has focused on supporting short-form summaries, a sentence or paragraph in length. In the broader area of human-LLM writing, prior work has looked into providing assistance for writing long-form creative texts (Chung et al., 2022; Goldfarb-Tarrant et al., 2019; Singh et al., 2022; Mirowski et al., 2023), but these tasks do not take a long-form text input. To the best of our knowledge, no work has developed and evaluated a mixed-initiative, LLM-powered tool for generating long-form summaries of scientific long documents, such as research-paper blog posts. Shamsabadi and D’Souza presented a prompt-based tool to support different research tasks including writing blog posts; however, the tool’s main function is to generate a personalized prompt for the user to then provide to an LLM, and the tool was not evaluated (Shamsabadi and D’Souza, 2024). Still, prior work has shown that LLM-powered tools can support planning, drafting, and revising different writing forms, such as creative stories (Goldfarb-Tarrant et al., 2019), short summaries (Zhang et al., 2023b), and short scientific texts (Long et al., 2023). Thus, in this paper we ask: can an LLM-powered tool provide assistance in dealing with the challenges of writing long-form summaries of scientific long documents?
As indicated by the many guidelines that discuss the content, formatting, and language of academic blog posts (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), there are multiple challenges specific to summarizing a research paper in a blog post. The first distinct challenge is related to the input being a long document. The writer must sift through the paper to determine what information to carry over to the blog post, without missing important details. The two other main challenges are related to the output being long-form and for science communication. The writer must organize key points in sections amenable to a scientific blog post. Also, the writer must ensure that their writing is scientifically accurate like the paper but more concise, easy to understand, and eye-catching than the paper.
@@
To help researchers write research-paper blog posts, we developed Papers-to-Posts, a mixed-initiative tool designed based on guidelines for academic blog post writing (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Mehlenbacher, 2019) as well as pilot study feedback. The tool follows our Plan-Draft-Revise workflow (Figure 1). In the Plan phase, the user may adjust the tool’s selection of key bullet points within an LLM-generated paper outline. Inspired by prior work on LLM-supported reverse outlining (Dang et al., 2022), Papers-to-Posts is, to the best of our knowledge, the first LLM-powered long-document summarization tool to support direct control of content selection through an LLM-generated document outline, maintaining the source’s narrative structure. In the Draft phase, the user may customize preset instructions (e.g., include a hook, make one paragraph) to the LLM for each blog post section. The LLM then uses the selected bullet points and adjusted instructions to generate text for the given section. In the Revise phase, the user may further refine text with LLM-powered macros that encapsulate frequent editing transformations for scientific blog post writing (e.g., to reduce jargon, to better capture attention).
@@
We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers. In these studies, we compared our tool to a strong baseline tool–an LLM-generated draft and access to free-form LLM prompting. In a within-subjects lab study (N=20 participants) in which researchers wrote blog posts for two papers (one per tool), we found that Papers-to-Posts led to significantly greater change to the provided draft within a fixed amount of time alongside significantly higher writer satisfaction with the final blog post, without significantly increasing cognitive load. To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37 blog posts, 26 participants). Again, participants using Papers-to-Posts made greater change to the provided drafts for a given amount of time and for a given number of writing actions taken. Through surveys, interviews, and logs from across the two studies, we also observed evidence that an interactive LLM-generated paper outline for planning writing as well as preset but customizable LLM instructions for drafting and revising writing contributed to this increased editing power. In discussing our results, we reflect on implications for other human-LLM tools for research-paper summarization and long-form scientific communication: avoid solely post-editing interactions, provide an interactive outline of the long document for planning the summary, and include preset but adjustable LLM instructions for generating and modifying the writing.
@@
  • Papers-to-Posts, a tool for writing research-paper blog posts that implements the Plan-Draft-Revise workflow (Figure 1 and Section 3), including providing an interactive LLM-generated paper outline for planning the blog post, section-specific and customizable LLM instructions for drafting the blog post, and macros for revising text style to suit the blog post.
  • Findings from a within-subjects lab study (N=20 participants) showing that Papers-to-Posts significantly increases researchers’ editing power (i.e., change in writing within a fixed amount of time) (Figure 5a and Section 5.1.1), while significantly increasing satisfaction with the output blog post (Figure 6a and Section 5.2.1) and without significantly impacting cognitive load.
  • Findings from a between-subjects deployment study (N=37 blog posts, 26 participants) demonstrating that Papers-to-Posts increases researchers’ editing power (i.e., change in writing within an amount of time or writing actions) (Figure 5b-c and Section 5.1.2), without decreasing satisfaction with the output blog post or increasing cognitive load.
@@ 2.1. Human-LLM Text Summarization
Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022). To address these issues, other works have looked into having a human and LLM work together to generate a summary (Lai et al., 2022; Moramarco et al., 2021). There are several works on human-LLM summarization of long documents (Zhang et al., 2011; Gehrmann et al., 2019) and some for scientific long documents in particular (Zarinbal et al., 2019; Bayatmakou et al., 2022; Zhang et al., 2023b; Cheng et al., 2023). Long documents provide a unique challenge for human-LLM summarization, as there is a great deal of information for the human to digest. As writing research-paper blog posts entails summarizing portions of a research paper, this work seeks to support human-LLM long-document summarization.
According to Cheng et al., there are five common methods of human-AI summarization: guiding AI output, selecting or rating AI output, post-editing AI-generated summaries, interactive editing, and writing with AI assistance (Cheng et al., 2022). Our tool Papers-to-Posts provides affordances related to guiding and post-editing AI output. Prior work in human-AI summarization has mainly focused on model guidance through various means (e.g., concepts/queries (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023; Wang et al., 2023b), typing (Xie et al., 2023), ratings (Gao et al., 2020; Bohn and Ling, 2021), gaze (Bohn and Ling, 2021)) or simply post-editing (e.g., (Lai et al., 2022; Moramarco et al., 2021)). A few works have supported summarization based on user-selected points from the source artifact (Bayatmakou et al., 2022; Yan et al., 2011; Zhang et al., 2011; Wang et al., 2023a, b), but the points available for selection are ranked or clustered by the system without the same narrative structure as the source. Another work proposed but did not implement or evaluate model guidance through in-context content selection (Gehrmann et al., 2019). Our approach allows model guidance directly grounded in the input source while maintaining the source narrative structure.
Providing more direct control of input content is useful when the writer has significant knowledge of the source and can use reminders of its contents when selecting what to include in the summary. We focus on summarizing one’s own content, unlike most prior work; the exception is work on summarizing medical notes, which used the post-editing technique (Moramarco et al., 2021; Cai et al., 2022). We take advantage of the writer’s knowledge of the source document by having them select content within the context of the source’s layout.
@@ 2.2. Outlines in Human-LLM Writing
With respect to planning blog post text, a unique aspect of our work is providing an interactive LLM-generated outline for source content selection. There are some works in human-LLM writing that have utilized outlines for writing, but they focused on outlining output content rather than input content. For example, Zhang et al. introduced the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a), and Lu et al. created a tool to support writing workplace social media posts that includes a feature to help generate an outline for the post (Lu et al., 2024). Most related to our work, Dang et al. introduced a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022). The summaries then become affordances to help writers revise their writing (e.g., restructuring the writing). Our work also generates a reverse outline, but it is an outline of the input paper for users to interactively guide blog post generation. For summarizing a long source document, an outline is particularly useful, as reviewing the entire source may require substantial time and effort.
2.3. Preset, Customizable LLM Instructions in Human-LLM Writing
For drafting blog post text, our tool presents LLM instructions preset to generate four unique research-paper blog post sections; the user may then adjust these section-specific instructions in multiple manners to draft text that better aligns with their needs. Prior works also break text into chunks to support drafting and revising writing. For example, Laban et al. present an editing interface that provides edit suggestions for relatively small chunks of writing (Laban et al., 2023), while Dhillon et al. provide next-sentence and next-paragraph suggestions to continue writing (Dhillon et al., 2024). There are also prior works on human-LLM writing that, like ours, allow users to provide customized instructions to LLMs for drafting text (e.g., (Lim et al., 2024; Coenen et al., 2021)). Our customizable LLM instructions for drafting text are specifically designed to support writing common sections of research-paper blog posts.
For revising blog post text, our tool provides preset macros as well as the ability to perform custom text modifications. Several existing tools (e.g., (Google, [n. d.]; NotionAI, [n. d.]; Lu et al., 2024; Coenen et al., 2021; Reza et al., 2024)) also allow users to modify their text with macros for actions such as adjusting the text’s length, simplicity, or professionalism, and some allow custom modifications. However, our tool’s macros are designed specifically for blog post writing.

### v3

# Papers-to-Posts: Supporting Detailed Long-Document Summarization with an Interactive LLM-Powered Source Outline

@@ Abstract.
Compressing long and technical documents (e.g., ¿10 pages) into shorter-form articles (e.g., ¡2 pages) is critical for communicating information to different audiences, for example, blog posts of scientific research paper or legal briefs of dense court proceedings. While large language models (LLMs) are powerful tools for condensing large amounts of text, current interfaces to these models lack support for understanding and controlling what content is included in a detailed summarizing article. Such capability is especially important for detail- and technical-oriented domains, in which tactical selection and coherent synthesis of key details is critical for effective communication to the target audience. For this, we present interactive reverse source outlines, a novel mechanism for controllable long-form summarization featuring outline bullet points with automatic point selections that the user can iteratively adjust to obtain an article with the desired content coverage. We implement this mechanism in Papers-to-Posts, a new LLM-powered system for authoring research-paper blog posts. Through a within-subjects lab study (n=20) and a between-subjects deployment study (n=37 blog posts, 26 participants), we compare Papers-to-Posts to a strong baseline tool that provides an LLM-generated draft and access to free-form prompting. Under time constraints, Papers-to-Posts significantly increases writer satisfaction with blog post quality, particularly with respect to content coverage. Furthermore, quantitative results showed an increase in editing power (change in text for an amount of time or writing actions) while using Papers-to-Posts, and qualitative results showed that participants found incorporating key research-paper insights in their blog posts easier while using Papers-to-Posts.
Figure 1. Interactive reverse source outlines in the Papers-to-Posts system. Users input a long-form source document (research paper) and the system provides (a) a WARM START for the summarization task by generating a reverse document outline. The system then produces a draft summary article (blog post) with sections based on outline bullet points (and associated source paragraphs) selected by the LLM. Given the draft, users can perform two main actions: (b) SUMMARY REVISION, where users adjust the system’s bullet point selection for the summary, triggering system re-generation of the summary; and (c) SECTION ADDITION, where users provide a header for a new desired section, based on which the system selects bullet points and generates a draft of the section content. Users may also edit the draft manually.
@@
The figure has three rows and two columns. Each row describes an action supported by the interactive reverse source outlines. The first column describes the action, and the second column illustrates the action. The first row, labeled ”a,” is titled ”Warm Start.” It says the following. ”User [in orange]: Provide source document. Tool [in blue]: Write source outline, select relevant outline points, and generate summary.” Meanwhile, the illustration shows a set of orange papers labeled ”research paper.” A blue arrow points from the research paper to an image (labeled ”interactive paper outline”) of an outline with sets of 2 or 3 bullet points, some checked off with a blue check-mark and some not. A blue arrow points from the outline to an image labeled ”blog post.” It contains four sections labeled ”introduction,” ”methods,” ”results,” and ”conclusion.” There are blue lines in all four section of the blog post. The second row, labeled ”b,” is titled ”Summary Revision.” It says the following. ”User [in orange]: Review and revise selected outline points. Tool [in blue]: Re-generate summary based on updated point selection.” The illustration shows the interactive paper outline updated so that an orange ”X” covers one of the blue checkmarks, and an orange checkmark is added for one of the bullet points previously unchecked. A blue arrow points from the outline to the blog post, which is updated to have some of its blue lines turned orange. The length of the conclusion section also changed. The third row, labeled ”c,” is titled ”Section Addition.” It says the following. ”User [in orange]: Add a new summary section header. Tool [in blue]: Select relevant outline points for new section based on provided header and generate section. In the illustration, the outline is updated to have no orange marks and blue checkmarks on different bullet points than previously. A blue arrow points from the outline to the blog post, which contains a new section added as the second section. It is titled ”Background” in orange and its lines are in blue. The other blog post sections’ lines representing writing are in gray.
@@ 1. Introduction
Condensing long and technical documents into shorter forms for different audiences is crucial for many applications, from blog posts about scientific papers to legal briefs about court proceedings. In science, communicating scientific ideas from research papers to broader audiences increases awareness and impact (Jucan and Jucan, 2014; Burns et al., 2003). Scientists often seek different ways to publicize their work, such as through tweets or blog posts (Gero et al., 2021; Long et al., 2023; Williams et al., 2022; Holmberg and Thelwall, 2014); indeed, tweeting about research papers has been shown to raise their citation count (Luc et al., 2021). While some tools have been created to help author short-form derivatives like tweets (Long et al., 2023), significant challenges remain for scientists seeking to create more detailed derivatives like blog posts. Summarizing long documents (10k+ words) in detailed articles (500-1000 words) takes considerable time and effort, even with LLM support. Detailed articles require the writer to identify large amounts of relevant content from the source document. Given a long document, identifying all the important information to include can be daunting. While LLMs have impressive abilities to synthesize information from a long piece of text, what the LLM chooses to leave out is unclear to the user, and quickly iterating on the selected content is difficult. In domains like science, law, and business, attention to detail is critical; a summary missing key content could lead to misrepresented work and misinformed readers.
While some prior work has investigated fully automatic summarization of long documents (Koh et al., 2022), a mixed-initiative approach allows users to have more control over their summaries, which is important in detail-oriented domains like scientific research. Prior work in human-AI text summarization has often focused on helping create short-form summaries around a paragraph in length, which are easier to control through selection of a few source-document sentences (Zhang et al., 2011; Wang et al., 2023). However, when writing longer summaries, selecting source content at the sentence level becomes unwieldy. Other work in human-AI text summarization has looked into other interactions that may support longer-form summaries, but these works involve interactions detached from the source narrative, like post-editing of an automatically generated summary (Lim et al., 2024; Lai et al., 2022; Moramarco et al., 2021) or selecting concepts to include in the summary (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023). Without sufficient context from the source document, making sure that important details are not missing in a longer summary is difficult.
Inspired by prior work on LLM-supported reverse outlining for general writing (Dang et al., 2022), we present interactive reverse source outlines, a new interaction mechanism for controlling summary content selection with an interactive LLM-generated source outline, which maintains the source’s narrative structure and simplifies the task of selecting relevant content (Figure 1). To the best of our knowledge, this is the first mixed-initiative approach for detailed and grounded summarization of long documents (Figure 2).
@@
We implement this mechanism in Papers-to-Posts, a mixed-initiative system that helps researchers write research-paper blog posts. The system follows a plan-draft-revise writing workflow (Flower and Hayes, 1981). In the planning phase, the user may review the system’s initial blog post draft and adjust the system’s selection of key bullet points within an LLM-generated paper outline. In the drafting phase, the user may customize preset instructions (e.g., include a hook, write one paragraph) to the LLM for each blog post section. The LLM then uses the selected bullet points and adjusted instructions to generate new text for the given section. In the revising phase, the user may further refine text with LLM-powered macros that encapsulate frequent editing transformations for scientific blog post writing (e.g., to reduce jargon, to better capture attention).
@@
We evaluated Papers-to-Posts through two studies in which participants wrote blog posts for their own papers. In these studies, we compared our tool to a strong baseline tool–an LLM-generated draft and access to free-form LLM prompting. In a within-subjects lab study (N=20 participants) in which researchers wrote blog posts for two papers (one per tool), we found that Papers-to-Posts led to significantly higher writer satisfaction with the final blog post, and participants were particularly more satisfied with the content covered in the final blog post. Furthermore, participants had higher confidence that all essential information was in the final blog post when using Papers-to-Posts. To study Papers-to-Posts in a more realistic setting, we also conducted an unmonitored, between-subjects deployment study (N=37 blog posts, 26 participants). Across both studies, participants found Papers-to-Posts more helpful. They found it easier to incorporate research-paper content in the blog post and to iterate on that content. Moreover, participants demonstrated increased editing power (change to the provided blog post draft for a given amount of time or for a given number of writing actions) without an increase in cognitive load.
@@
  • Interactive reverse source outlines, a novel mechanism for more user control in human-LLM detailed summarization of long documents, consisting of an LLM-generated reverse outline of the source document with pre-selected bullet points for an initial draft of the summary, with which the writer iteratively interacts to control content selection
  • Papers-to-Posts, a tool that implements this mechanism for the purpose of writing research-paper blog posts, a common long-document summarization task that requires attention to detail
  • Findings from a within-subjects lab study (N=20 participants) and a between-subjects deployment study (N=26 participants) showing that Papers-to-Posts
    • Under time constraints, significantly increases writers’ satisfaction with their summaries’ quality (primarily with respect to content coverage)
    • Is considered more helpful by writers, making incorporating source content in a detailed summary easier and iterating on the summary easier, with increased editing power (change in writing within an amount of time or writing actions) and no increase in cognitive load
@@ 2.1. Human-LLM Text Summarization
Although there is much work on automatic text summarization (Koh et al., 2022; El-Kassas et al., 2021), issues remain with the output summaries, such as inaccuracies and style misalignment (Lu et al., 2023; Koh et al., 2022). To address these issues, other works have looked into supporting humans in working together with LLMs to generate a summary (Cheng et al., 2022). These tools may be divided into groups in terms of 1) the summary length that they support and 2) how grounded the interaction is in the context of the source document (Figure 2).
Some prior work in human-LLM text summarization supports an interaction in which the human selects sentences extracted directly from the source (Wang et al., 2023; Zhang et al., 2011) (Figure 2, top left). For a short summary, identifying the few key sentences from the source to include may be manageable, but for a detailed summary, this would require considerable time and effort. One work enables more detailed summaries by generating expanded text based on the user selecting a few source sentences (Bayatmakou et al., 2022) (Figure 2, middle right), but what source content has been included versus neglected then becomes unclear. Other prior work provides users with automatically generated summaries that they can adjust through interactive editing (e.g., via ratings, gaze) (Bohn and Ling, 2021; Gao et al., 2020; Xie et al., 2023; Yan et al., 2011) or post-editing (Moramarco et al., 2021; Cai et al., 2022; Lai et al., 2022) (Figure 2, bottom right). While these interactions are flexible in terms of the length of summary generated, they do not support users in understanding what source content has or has not been represented. Yet another group of prior works strikes a balance between generating longer summaries and grounding them in the source document, by having the user select concepts or keywords to include in the summary (Zhang et al., 2023b; Bayatmakou et al., 2022; Avinesh and Meyer, 2017; Avinesh et al., 2018; Ding et al., 2023) (Figure 2, center). Concepts are easier to sort through than full sentences, which makes generating a longer summary more feasible, and selecting concepts provides users some sense of what source content is utilized overall. Nevertheless, concepts are high-level and abstract, limiting their utility for generating detailed summaries and easily identifying what source details are and are not included. Our novel interactive reverse source outlines, implemented in Papers-to-Posts (Figure 2, top right), support both grounded and long-form summarization by allowing users to select outline bullet points directly tied to the source document in order to generate a detailed summary.
Figure 2. Comparing interactive reverse source outlines in Papers-to-Posts (blue, upper right) with other affordances for control of LLM summarization, in terms of 1) how grounded the affordance is in the source narrative and 2) the length of the output summary that is supported. Papers-to-Posts is the first LLM-powered tool to provide writers with highly grounded control over a detailed summary.
[…]
A scatter plot titled ”Affordances for Control of LLM Text Summarization.” The x-axis is an arrow pointing right, titled ”output summary length.” The left side of the arrow is labeled ”paragraph” and the right side is labeled ”detailed article.” The y-axis is an arrowing pointing upward, titled ”grounding in source narrative.” The bottom of the arrow is labeled ”ungrounded,” and the top is labeled ”grounded.” The scatterplot has five dots: 1) a dot in the upper left labeled ”select extracted sentences [44,50]”; 2) a dot near the center labeled ”select concepts [2-4,15,49]”; 3) a dot in the bottom right labeled ”interactively edit [6,18,46,47] or post-edit [10,30,39] generated text”; 4) a dot around the middle right labeled ”expand on selected extracted sentences [4]”; and 5) a dot in the upper right labeled ”select points from generated outline (PAPERS-TO-POSTS).”
@@ 2.2. Outlines in Human-LLM Writing
There are some works in human-LLM writing that have utilized LLM-generated outlines for writing. Some focus on outlining output content, rather than outlining input content for summarization (Zhang et al., 2023a; Lu et al., 2024). For instance, Zhang et al. introduce the tool VISAR, a human-AI argumentative writing assistant that allows users to iterate on their draft with adjustments to a visual outline of the draft (Zhang et al., 2023a). Other works focus on outlining input content (as we do), but the outline is not utilized to synthesize a detailed summary of the input (Petridis et al., 2023; Yu et al., 2024). As an example, AngleKindling generates a few main points from an input press release as a quick overview for journalists to then ideate about potential angles for reporting on the press release (Petridis et al., 2023). However, there is no support for interacting with the main points to generate a summary of the press release.
Most related to our work, Dang et al. present a text editor that generates summaries of each paragraph written in the editor to comprise a reverse outline of the writing (Dang et al., 2022). The summaries then become affordances to help writers think about how to revise their writing. Our work also generates a reverse outline, but it is an outline of the input document for users to interactively guide summary generation. For summarizing a long source document, an outline is particularly useful, as reviewing the entire source itself requires substantial time and effort.